# Batched Bayesian optimization with correlated candidate uncertainties

### v1

@@ 4.3 Application to the design of organic electronics
We next demonstrate qPO on the pursuit of molecules from the QM9 dataset of 133K compounds that maximize the DFT-calculated HOMO-LUMO gap (Ramakrishnan et al., 2014; Ruddigkeit et al., 2012). We begin each search with a randomized initial batch of 100 compounds and select 100 more at each of 20 subsequent iterations. Optimization performance according to top-$k$ metrics is shown in Figure 4. As in the previous study, qPO performs as well as or better than all baselines. In this case, qEI is the most competitive baseline, performing comparably with or very slightly worse than qPO across all evaluation metrics. We observe the greatest improvement to top-$k$ optimization performance when considering small values of $k$ (e.g., Figure 4A and 4D). This result aligns with qPO’s primary focus on identifying the true global optimum.
Figure 4: Optimization performance on the exploration of the 133K QM9 dataset for compounds with large HOMO-LUMO gaps. (A-C) Retrieval of the true top 0.01%, 0.1%, and 1%, representing 14, 134, and 1,339 top-performing compounds, respectively. (D-F) Average oracle values of the top 10, 50, and 100 acquired compounds. Shaded regions denote $\pm$ one standard error of the mean across ten runs (A.5, Tables 2 and 3). qPO, multipoint probability of optimality (ours); pTS, parallel Thompson sampling; qEI, multipoint expected improvement; UCB, upper confidence bound.
@@ 5 Conclusion
We have proposed an exploitative batch acquisition function (qPO) for discrete Bayesian optimization that maximizes the likelihood that the batch contains the true optimum. qPO is equal to the sum over individual acquisition scores and therefore circumvents the combinatorial challenge of optimizing a batch-level acquisition score. We explain how the treatment of model covariance implicitly captures diversity and how it differentiates qPO from parallel Thompson sampling in subtle but meaningful ways. Empirically, our method efficiently identifies an equivalent or higher percentage of top-performing candidates when compared to batched BO alternatives. The most notable improvement to top-$k$ metrics is observed for smaller values of $k$, consistent with the acquisition strategy’s goal of acquiring the true global optimum.
The proposed batch acquisition strategy has some notable limitations. First, qPO cannot be computed analytically (in an exact sense), necessitating a Monte Carlo estimate. Alternative methods for estimating the probability of a candidate’s optimality, such as expectation propagation, may reduce the computational cost of implementing qPO. Second, qPO may fail to select diverse batches in the presence of high observation noise, which will reduce the relative impact of covariance on qPO scores222 For example, if $\lambda I$ were added to the covariance matrix in Eq. 12, $(x_{1},x_{2})$ would be the optimal batch for sufficiently large $\lambda$. . Theoretical analysis of the exploration-exploitation trade-off in the finite iteration setting may uncover potential failure modes of qPO and further support the empirical results observed here.
Acknowledgments
This work was supported by the Office of Naval Research under Grant No. N00014-21-1-2195 and the Machine Learning for Pharmaceutical Discovery and Synthesis consortium. JF was supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064. MM acknowledges support from funding from the National Institutes of Health under Grant. No. 605 T32GM087237. JMHL acknowledges support from a Turing AI Fellowship under grant EP/V023756/1.
Reproducibility Statement
All code required to generate the results shown in this work, including installation instructions, the datasets explored in Section 4, and code to run BO with qPO, can be found at https://github.com/jenna-fromer/qPO.
Appendix A Appendix
A.1 The non-unique optimum case
In some cases, the mutual exclusivity of events $\{x^{*}=x_{i}\}_{x_{i}\in\mathcal{X}}$ may not be assumed during the estimation of qPO following Eq. 4.
In such scenarios, we will not arrive at the result in Eq. 5 and must instead optimize the acquisition function by constructing batches myopically (i.e., one-by-one):
@@ A.1 The non-unique optimum case
The methods described in Section 2.3 are also applicable to optimizing this batch acquisition function.
Note that, even when the events $\{x^{*}=x_{i}\}_{x_{i}\in\mathcal{X}}$ are not guaranteed to be mutually exclusive, observation noise in the surrogate model may allow us to assume that the events $\{\hat{f}(x^{*})=\hat{f}(x_{i})\}_{x_{i}\in\mathcal{X}}$ are mutually exclusive. In this case, we can apply the acquisition strategy in Eq. 5 because all predictions of the events $\{x^{*}=x_{i}\}_{x_{i}\in\mathcal{X}}$ will be mutually exclusive, and estimations of the conditional probabilities in Eq. 13 will in practice be equivalent to the corresponding unconditional probabilities.
A.2 Proof of bound for Eq. 11
If
$\Pr(x^{*}=x_{i})\geq 1-\delta^{\frac{1}{M}}\ ,\$
then the probability of never observing $\left(x^{*}=x_{i}\right)$ in $M$ i.i.d. Monte Carlo samples is less than $\delta$ (in equation 11).
Proof.
The event of $x_{i}=x^{*}$ is binary and therefore has a Bernoulli distribution. For a Bernoulli distribution with expectation $q$, the probability of observing only negative events for $M$ i.i.d. samples is $\left(1-q\right)^{M}$ and is decreasing with $q$. Re-arranging this gives the bound above. ∎
This result shows that, with modest $M$, inputs $x$ which have a reasonable probability of being the optimum are very unlikely to never be observed as the optimum in some sample. For example, any candidate with at least a 0.7% chance of being the optimum has less than a 0.1% chance of never appearing as the optimum with $M=10^{3}$ Monte Carlo samples.
We can use Hoeffding’s inequality to get a more precise confidence interval:
Lemma 1.
Let $p=\Pr\left(x_{i}=x^{*}\right)$ and let $\hat{p}$ be a Monte Carlo estimate for $p$ with $M$ independent samples. Let $\epsilon=\sqrt{\frac{\log{2/\alpha}}{2M}}$. Then the set $C=\left(\hat{p}-\epsilon,\hat{p}+\epsilon\right)$ is a $1-\alpha$ confidence interval for every $p$ satisfying
$\Pr\left(p\in C\right)\geq 1-\alpha\ .$
Proof.
From Hoeffding’s inequality we have
$\Pr\left(\left|p-\hat{p}\right|\right)\leq 2\exp{\left(-2M\epsilon^{2}\right)}\ .$
Solving $\alpha=2e^{-2n\epsilon^{2}}$ gives the desired result. ∎
Setting $\alpha=0.1\%$ to give a 99.9% confidence interval, if a particular candidate $x_{i}$ is never observed to be the optimum in $M=10^{3}$ samples ($\hat{p}=0$), then $p\in\left(-0.062,0.062\right)$ with 99.9% probability.
A.3 Recasting acquisition scores as orthant probabilities
The acquisition score in Eq. 7 can be defined as an orthant probability through a change of variables. This enables the estimation of $\Pr(x^{*}=x_{i})$ without posterior sampling. However, efficient calculation of orthant probabilities is not possible for arbitrary probability distributions. Here, we choose to require the posterior to be a multivariate Gaussian (or approximated as such). The likelihood of candidate $x_{i}$ being optimal is equivalent to the probability that $y_{i}$ is greater than all $y_{j,j\neq i}$:
$\Pr(x^{*}=x_{i})=\Pr(y_{i}>y_{1},y_{i}>y_{2},...,y_{i}>y_{N})$ (14)
Following the approach of Azimi et al. (2010), we denote the difference between $y_{i}$ and $y_{j}$ as $z_{j}^{i}$ and the vector of differences for candidate $i$ as $\mathbf{z}^{i}\in\mathbb{R}^{N-1}$. Using the transformation matrix $\textbf{A}_{i}\in\mathbb{R}^{(N-1)\times N}$ as defined by Azimi et al. (2010) and Azimi (2012), we may define:
$\mathbf{z}^{i}\sim\mathcal{N}\left(\textbf{A}\bm{\mu}_{\textbf{x}},\textbf{A}\bm{\Sigma}_{\textbf{x}}\textbf{A}^{T}\right)$ (15)
where $\bm{\mu}_{\textbf{x}}$ and $\bm{\Sigma}_{\textbf{x}}$ parameterize the surrogate model posterior for candidates x. The expression in Eq. 14 is equal to the orthant probability of $\mathbf{z}^{i}$, i.e., the probability that all elements of $\mathbf{z}^{i}$ are greater than 0. Azimi et al. (2010) apply a whitening transformation to the distribution defining $\mathbf{z}^{i}$, decorrelating all entries and enabling approximation of the orthant probability. Alternative methods to estimate high-dimensional orthant probabilities (Miwa et al., 2003; Craig, 2008; Ridgway, 2016) may also be applied here.
A.4 Experimental details
All code required to reproduce the results for this work can be found at https://github.com/jenna-fromer/qPO.
A.4.1 Implementation of qPO acquisition strategy
@@
A.4.2 Bayesian optimization
We apply our batch acquisition strategy to the model-guided exploration of two discrete chemical libraries (Section 4). The first case study explores a library of 39,312 compounds for putative antibiotics that minimize the growth of Staphylococcus aureus (Wong et al., 2024). We randomly initialize our search with 50 compounds and their growth inhibition values and acquire batches of 50 compounds for 10 subsequent iterations. In the second case, we explore the QM9 dataset for compounds that maximize the HOMO-LUMO gap (Ramakrishnan et al., 2014; Ruddigkeit et al., 2012). Here, we use an initial batch size of 100 compounds and acquire 100 more for 20 subsequent iterations. For both case studies, ten runs were performed for each acquisition method with distinct random seeds that govern the initial batch.
Our surrogate model is a Gaussian process with a Tanimoto kernel (Tanimoto, 1958) and a constant mean, a common surrogate model architecture for molecular BO (Tripp et al., 2021; García-Ortegón et al., 2022; Gao et al., 2022). Compounds are featurized as 2048-length count Morgan fingerprints using rdkit (Landrum, 2024). At each iteration, the Gaussian process hyperparameters—mean, covariance scale, and likelihood noise—are optimized by maximizing the marginal log likelihood of the model over all previously acquired training data.
A.4.3 Acquisition functions
Selecting batches with qEI, pTS, and qPO can be computationally expensive for large design spaces due to the $O(N^{3})$ complexity of Cholesky decomposition for $N$ candidates. To reduce the computational cost of these acquisition functions, we first filter the set of $N$ candidates to the top $10{,}000$ based on predicted mean and subsequently apply the appropriate acquisition function to select from these $10{,}000$ candidates. This step is not necessary for the implementation of these acquisition functions but facilitates their use under computational resource limitations.
@@ A.4.3 Acquisition functions
Upper confidence bound
@@
Parallel Thompson sampling
We followed the implementation of Hernández-Lobato et al. (2017), summarized in Algorithm 2. For each of $b$ posterior samples, the candidate $x_{i}$ which optimizes the objective and is not already in the batch is selected.
Multipoint expected improvement
We implement qEI using the qLogExpectedImprovement (Ament et al., 2023) and optimize_acqf_discrete functions in botorch (Balandat et al., 2020). qLogExpectedImprovement evaluates qEI using Monte Carlo sampling. optimize_acqf_discrete sequentially selects points and appends their corresponding $x$ values to the Gaussian process input for subsequent Monte Carlo samples, making this a hallucinating approach.
@@
A.4.4 Analyzing batch diversity

### v2

@@ 4.3 Application to the design of organic electronics
We next demonstrate qPO on the pursuit of molecules from the QM9 dataset of 133K compounds that maximize the DFT-calculated HOMO-LUMO gap 54, 55. We begin each search with a randomized initial batch of 100 compounds and select 100 more at each of 20 subsequent iterations. Optimization performance according to top-$k$ metrics and total run time is shown in Table 2, and top 0.01% retrieval of qPO and select baselines is plotted for all iterations in Figure 3B. In contrast to the previous demonstration, BUCB appears to be the most competitive method. Nonetheless, as in the previous study, qPO is among the top-performing methods. We observe the greatest improvement to top-$k$ optimization performance when considering small values of $k$ (Figure 3), aligning with qPO’s primary focus of identifying the true global optimum. The computational expense of qPO in this experiment, in terms of total run time, is comparable to that required by other batch acquisition strategies including BUCB, TS-RSR, and qPI.
Method Iteration Top 10 Average ($\uparrow$) Top 100 Average ($\uparrow$) Fraction Top 0.01% ($\uparrow$) Fraction Top 1% ($\uparrow$) Wall Time (s) ($\downarrow$)
BUCB 10 0.45 $\pm$ 0.01 0.39 $\pm$ 0.00 0.80 $\pm$ 0.07 0.31 $\pm$ 0.01 1401 $\pm$ 6
DPP-TS 10 0.42 $\pm$ 0.01 0.38 $\pm$ 0.00 0.41 $\pm$ 0.12 0.30 $\pm$ 0.00 02902 $\pm$ 36
Greedy 10 0.40 $\pm$ 0.01 0.38 $\pm$ 0.00 0.09 $\pm$ 0.09 0.36 $\pm$ 0.02 1020 $\pm$ 0
TS-RSR 10 0.45 $\pm$ 0.01 0.39 $\pm$ 0.00 0.71 $\pm$ 0.12 0.36 $\pm$ 0.01 01834 $\pm$ 13
UCB 10 0.40 $\pm$ 0.01 0.38 $\pm$ 0.00 0.10 $\pm$ 0.09 0.37 $\pm$ 0.01 1026 $\pm$ 8
pTS 10 0.40 $\pm$ 0.01 0.38 $\pm$ 0.00 0.15 $\pm$ 0.10 0.23 $\pm$ 0.01 01855 $\pm$ 13
qEI 10 0.43 $\pm$ 0.01 0.38 $\pm$ 0.00 0.46 $\pm$ 0.14 0.28 $\pm$ 0.02 1974 $\pm$ 3
qPI 10 0.44 $\pm$ 0.01 0.39 $\pm$ 0.00 0.60 $\pm$ 0.12 0.27 $\pm$ 0.01 1385 $\pm$ 6
qPO 10 0.42 $\pm$ 0.01 0.38 $\pm$ 0.00 0.40 $\pm$ 0.12 0.35 $\pm$ 0.01 1883 $\pm$ 9
random10k 10 0.38 $\pm$ 0.00 0.36 $\pm$ 0.00 0.01 $\pm$ 0.01 0.07 $\pm$ 0.00 1020 $\pm$ 2
BUCB 20 0.47 $\pm$ 0.00 0.40 $\pm$ 0.00 1.00 $\pm$ 0.00 0.54 $\pm$ 0.00 04840 $\pm$ 12
DPP-TS 20 0.46 $\pm$ 0.00 0.40 $\pm$ 0.00 0.94 $\pm$ 0.01 0.59 $\pm$ 0.01 05728 $\pm$ 43
Greedy 20 0.45 $\pm$ 0.01 0.39 $\pm$ 0.00 0.73 $\pm$ 0.11 0.65 $\pm$ 0.01 1046 $\pm$ 1
TS-RSR 20 0.47 $\pm$ 0.00 0.40 $\pm$ 0.00 1.00 $\pm$ 0.00 0.63 $\pm$ 0.00 03687 $\pm$ 18
UCB 20 0.46 $\pm$ 0.00 0.39 $\pm$ 0.00 0.93 $\pm$ 0.00 0.66 $\pm$ 0.01 1051 $\pm$ 8
pTS 20 0.46 $\pm$ 0.00 0.39 $\pm$ 0.00 0.93 $\pm$ 0.02 0.50 $\pm$ 0.01 03715 $\pm$ 19
qEI 20 0.46 $\pm$ 0.00 0.40 $\pm$ 0.00 0.99 $\pm$ 0.01 0.59 $\pm$ 0.01 6016 $\pm$ 9
qPI 20 0.46 $\pm$ 0.00 0.39 $\pm$ 0.00 0.95 $\pm$ 0.03 0.43 $\pm$ 0.01 04850 $\pm$ 12
qPO 20 0.47 $\pm$ 0.00 0.40 $\pm$ 0.00 1.00 $\pm$ 0.00 0.63 $\pm$ 0.01 03774 $\pm$ 19
random10k 20 0.39 $\pm$ 0.00 0.37 $\pm$ 0.00 0.04 $\pm$ 0.02 0.15 $\pm$ 0.00 1044 $\pm$ 2
Table 2: Optimization performance on the exploration of the 133K QM9 dataset for compounds with large HOMO-LUMO gaps at an intermediate and final iteration. We first report the average oracle value of the top 10 and 100 acquired designs. Retrieval of the true top 0.01% and 1% represents 14 and 1,339 top-performing compounds, respectively. We also report the average wall time for each method. All values denote average metrics $\pm$ one standard error of the mean across ten runs. Values with the best mean by each metric are bolded, and remaining values whose mean are within one standard error of the mean of the best performer are underlined.
@@ 5 Conclusion
We have proposed a batch acquisition function (qPO) for discrete Bayesian optimization, motivated by exploitation, that maximizes the likelihood that the batch contains the true optimum. qPO is equal to the sum over individual acquisition scores and therefore circumvents the combinatorial challenge of optimizing a batch-level acquisition score, differentiating it from alternative batch acquisition strategies like multipoint expected improvement and multipoint probability of improvement. We explain how the treatment of model covariance implicitly captures diversity and how it differentiates qPO from parallel Thompson sampling in subtle but meaningful ways. When compared to state-of-the-art baselines in batched Bayesian optimization in two empirical settings, qPO is among the top-performing methods, although no single strategy outperforms all others across all metrics. The optimal acquisition strategy in a molecular Bayesian optimization depends on both the dataset and the prioritized optimization metric(s), and our results demonstrate that qPO complements existing strategies and may be an optimal choice under certain circumstances.
The proposed batch acquisition strategy has some notable limitations. First, qPO cannot be computed analytically (in an exact sense), necessitating a Monte Carlo estimate. Alternative methods for estimating the probability of a candidate’s optimality, such as expectation propagation, may reduce the computational cost of implementing qPO. Second, qPO may fail to select diverse batches in the presence of high observation noise, which will reduce the relative impact of covariance on qPO scores. Theoretical analysis of the exploration-exploitation trade-off in the finite iteration setting may uncover potential failure modes of qPO and further support the empirical results observed here.
6 Methods
@@
6.1 Implementation of qPO acquisition strategy
@@
6.2 Bayesian optimization
We apply our batch acquisition strategy to the model-guided exploration of two discrete chemical libraries (Section 4). The first case study explores a library of 39,312 compounds for putative antibiotics that minimize the growth of Staphylococcus aureus 52. We randomly initialize our search with 50 compounds and their growth inhibition values and acquire batches of 50 compounds for 10 subsequent iterations. In the second case, we explore the QM9 dataset for compounds that maximize the HOMO-LUMO gap 54, 55. Here, we use an initial batch size of 100 compounds and acquire 100 more for 20 subsequent iterations. For both case studies, ten runs were performed for each acquisition method with distinct random seeds that govern the initial batch.
Our surrogate model is a Gaussian process with a Tanimoto kernel 56 and a constant mean, a common surrogate model architecture for molecular BO 57, 58, 59. Compounds are featurized as 2048-length count Morgan fingerprints using rdkit 60. At each iteration, the Gaussian process hyperparameters—mean, covariance scale, and likelihood noise—are optimized by maximizing the marginal log likelihood of the model over all previously acquired training data.
6.3 Acquisition functions
Selecting batches with qEI, pTS, and qPO can be computationally expensive for large design spaces due to the $O(N^{3})$ complexity of Cholesky decomposition for $N$ candidates. To reduce the computational cost of these acquisition functions, we first filter the set of $N$ candidates to the top $10{,}000$ based on predicted mean and subsequently apply the appropriate acquisition function to select from these $10{,}000$ candidates. This step is not necessary for the implementation of these acquisition functions but facilitates their use under computational resource limitations. A similar pre-filtering method has been applied by Moss et al. 49 in a molecular discovery setting to reduce computational cost.
@@ 6.3 Acquisition functions
Upper confidence bound (UCB)
@@
Batch upper confidence bound (BUCB)
We implement BUCB using the
qUpperConfidenceBound 50 and optimize_acqf_discrete functions in botorch 32.
optimize_acqf_discrete sequentially selects points and appends their corresponding $x$ values to the Gaussian process input for subsequent Monte Carlo samples, making this a hallucinating approach. Following Wilson et al. 50, we set $\beta=\sqrt{3}$.
Parallel Thompson sampling (pTS)
We follow the implementation of Hernández-Lobato et al. 24, summarized in Algorithm 2. For each of $b$ posterior samples, the candidate $x_{i}$ which optimizes the objective and is not already in the batch is selected.
Multipoint expected improvement (qEI)
We implement qEI using the
qLogExpectedImprovement 61 and optimize_acqf_discrete functions in botorch 32.
qLogExpectedImprovement evaluates qEI using Monte Carlo sampling. optimize_acqf_discrete sequentially selects points and appends their corresponding $x$ values to the Gaussian process input for subsequent Monte Carlo samples, making this a hallucinating approach.
Multipoint probability of improvement (qPI)
We implement qEI using the
qProbabilityOfImprovement and optimize_acqf_discrete functions in botorch 32.
qProbabilityOfImprovement evaluates qPI using Monte Carlo sampling.
optimize_acqf_discrete sequentially selects points and appends their corresponding $x$ values to the Gaussian process input for subsequent Monte Carlo samples, making this a hallucinating approach.
Determinantal point process Thompson sampling (DPP-TS)
We follow the iterative batch construction strategy in Algorithm 1 described by Nava et al. 27. We define the kernel as the Tanimoto similarity between 2048-length Morgan fingerprints. DPP-TS is an iterative batch construction policy; we allow for $1{,}000$ iterations of DPP-TS batch design in each BO iteration.
Thompson sampling with regret to sigma ratio (TS-RSR)
We follow Algorithm 1 in Ren and Li 26 for the implementation of TS-RSR. As in pTS, we obtain $b$ posterior samples. The design which optimizes the regret to sigma ratio in each sample is acquired.
General-purpose Information-Based Bayesian OptimisatioN (GIBBON)
We implement GIBBON 49 using the qLowerBoundMaxValueEntropy and optimize_acqf_discrete functions in botorch 32.
@@
6.4 Wall time
Tables 1 and 2 report the average wall time required by each strategy. Each run was performed on a single node with 48 CPUs and 2 GPUs, although not all resources were used by all strategies. Gaussian process surrogate models were trained on a single GPU. Acquisition functions implemented in botorch 32 (BUCB, qEI, qPI, GIBBON) leverage GPU acceleration for Monte Carlo sampling. Remaining acquisition functions (qPO, greedy, UCB, pTS, DPP-TS, TS-RSR, random10k) were evaluated using CPUs only.
6.5 Analyzing batch diversity